find the coordinates of the midpoint of HX given that H(-1,3) and X(7,-1)
Answer by Alan3354(69443)

find the coordinates of the midpoint of HX given that H(-1,3) and X(7,-1)
-----------
Find the average of x & y.
x = (-1 + 7)/2 = 3
y = (3 - 1)/2 = 1
--> (3,1)